# Formalism 25

### v2

@@ 1. Formalism and finitism
Abraham Robinson’s intellectual profile continues to inspire passions fifty years after his passing. A number of publications over the past few decades have analyzed his philosophical position, including [Benis-Sinaceur 1988, p. 603], [Gaifman 2012], [Sherry and Katz 2012], [Sanders 2020], [Weir 2024], and most recently [Erhardt 2024]. A majority of commentators take it for granted that Robinson’s mature position is best described as mathematical Formalism. However, Erhardt reads Robinson as a finitist (see Section 1.2 for a discussion of distinct meanings of the term) who rejects infinite totalities.
@@ 1.1. Meaning and reference
In Formalism 64, Robinson’s most fully developed philosophical paper, he writes (Robinson 1964/1979b, pp. 230–231, italics in the original), “[…] the notion of a particular class of five elements, e.g. of five particular chairs, presents itself to my mind as clearly as the notion of a single individual […]. By contrast, I feel quite unable to grasp the idea of an actual infinite totality. To me there appears to exist an unbridgeable gulf between sets or structures of one, or two, or five elements, on one hand, and infinite structures on the other hand […]” [^2] Robinson as quoted in [Erhardt 2024, p. 2]; emphasis on “unable to grasp” added.
@@ 1.2. Gaifman’s reading of Robinson
Erhardt thanks Gaifman for encouraging him to read Robinson’s text Formalism 64 [Erhardt 2024, p. 17]. Gaifman may have been the original source of a characterisation of Robinson as a finitist. Indeed, Gaifman claims in 2012:
@@ 1.3. Standard model
There is, however, another alternative: more strongly committing to finitism. This is the path taken by various constructivists, who rather than play the uninterpretable game of symbols choose, in a philosophically principled manner, to adopt a weaker form of mathematics. [Erhardt 2024, p. 15; emphasis added]
@@ 1.4. Robinson’s passage in context
As noted in Section 1.1, Robinson clarified his adjective ‘meaningless’ in his 1975 text where he stated that what he has in mind is an absence of a reference; namely, the term infinite totality does not refer to anything in either the real or a Platonic realm. Already in Formalism 64 he emphasized that he rejected the idea that infinite totalities exist “either really or ideally” [Robinson 1965, p. 230]. Here ‘really’ refers to the physical realm, and ‘ideally’ to a Platonic realm, as he states explicitly in the passage quoted above. It emerges that Erhardt’s presentation of Robinson’s position, as when Erhardt claims that it “places Robinson at odds with typical mathematical practice” [Erhardt 2024, p. 3], conflates the generic meaning of terms such as to grasp and meaningless, with the precise technical meaning attributed to such terms by Robinson.
@@ 1.5. Fermat’s Last Theorem: is the proof legitimate?
Though Fermat’s Last Theorem is a general result proven by illegitimate methods and that purports to say something about all natural numbers—constituting an illicit reference to actual infinity—one can derive from it a potentially infinite number of legitimate, material claims. Yet the fact that the proof is illegitimate on Robinson’s view is crucial to our appraisal of his view. [Erhardt 2024, p. 12; emphasis added]
@@
In opposition to nominalists (and a fortiori, fictionalists), Robinson is a devout realist with regard to each individual natural number. This is evidenced by his commitment to potential infinity, …777[Erhardt 2024, note 22]; emphasis on “devout realist” added.
@@
With regard to the issue of nominalism and fictionalism mentioned by Erhardt above, we note a pecularity of Erhardt’s approach. Erhardt is keen to cast Robinson as an instrumentalist, rather than one who regards infinite classes as useful fictions. Erhardt dismisses the suggestion that Robinson was a fictionalist, in spite of acknowledging that Leibniz, an avowed fictionalist, “had a profound effect” on Robinson [Erhardt 2024, p. 8 and note 22]. The rationale for such a dismissal is most peculiar: Erhardt considers that it would be anachronistic for Robinson, in 1964, to see himself as a fictionalist since Field’s ‘introduction of fictionalism’ occurred only in 1980 (see [Field 1980]). Erhardt is right to note that Field’s fictionalism, a form of nominalism, conflicts with Robinson’s repeated rejection of nominalism. But such a conflict arises only if fictionalism is assumed to be necessarily a nominalist proposal. Leibniz’s fictionalism offers an alternative to nominalism. Leibniz treated infinitesimals, along with negative and imaginary numbers, as well-founded fictions; see [Sherry and Katz 2012]. Although he sometimes suggested that infinitesimals are eliminable in the manner of Archimedean exhaustion arguments, he made no such suggestion for negatives and imaginaries. In all three cases, it is the contribution to systematicity that establishes the mantle ‘well-founded fiction’, rather than any sort of nominalist reduction.
@@ 1.6. Symptomatic keyword list and imaginary tensions
The foundational position he inherited from David Hilbert undermines not only the use of nonstandard analysis, but also Robinson’s considerable corpus of pre-logic contributions [^8] Erhardt’s wording here constitutes a historical error. He makes it appear as though Robinson’s work in aeronautics, such as his 1956 book Wing Theory [Robinson and Laurmann 1956], preceded his work in logic. This is incorrect. For example, Robinson’s dissertation The metamathematics of algebraic systems [Robinson 1949], advised by Paul Dienes, dates from 1949; see https://www.genealogy.math.ndsu.nodak.edu/id.php?id=15886. See also Dauben [Dauben 1995, p. 157]. to the field in such diverse areas as differential equations and aeronautics. This tension emerges from Robinson’s disbelief in the existence of infinite totalities … [Erhardt 2024, Astract; emphasis added]
@@ 2. Twin prime conjecture and types of infinite totalities
[E]ven basic conjectures of number theory, such as the twin prime conjecture, presuppose the existence of infinite totalities. [Erhardt 2024, p. 2]
However, such a claim involves a conflation of distinct meanings associated with the term infinite totality; namely, of different levels of language or theory. The twin prime conjecture (TPC) can be formulated in Peano Arithmetic (PA). This can be done, for example, by the formula on page 13, line 4 in [Erhardt 2024], which we will express as follows:
@@ 3.1. A logical fallacy
While Robinson has no difficulty grasping the possibility of arbitrarily large finite sets, which motivates his acceptance of the natural numbers taken individually, he reports that it is simply a psychological fact that he cannot grasp infinitary objects, obviating the possibility of finding common ground with disputants.101010[Erhardt 2024, p. 2]; emphasis on ‘grasping’ in the original; emphasis on ‘reports’ and ‘psychological’ added.
@@ 3.2. Robinson’s letter to Gödel
Erhardt’s attribution of Alice’s position to Robinson constitutes a logical fallacy [^13] Robinson is not the only mathematician whose position is misrepresented by Erhardt. Erhardt claims that Edward Nelson “believe[d] there is a largest number” [Erhardt 2024, p. 4]. To support his claim, Erhardt cites four texts by Nelson. We examined all four, including Nelson’s publication [Nelson 2011], but found no evidence of such a belief as claimed by Erhardt. Erhardt’s claim in [Erhardt 2024, note 7] that Nelson “believed Peano Arithmetic was inconsistent” is similarly inaccurate. Rather, Nelson suspected that PA was inconsistent. When he thought he had a proof he obviously thought it was inconsistent, but when Tao found an error in the proof, Nelson immediately accepted it. and does not inspire confidence in his analysis.
@@ 3.3. Verbal excesses from Kreisel to Erhardt
Robinson has already admitted that such totalities are meaningless, and without an argument defending the use of these languages, formalism becomes an inconsistent project [Erhardt 2024, p. 6; emphasis added].
@@ 3.4. Whose game was it anyway?
“He explains non-finitary mathematics as a collection of ‘uninterpretable games with symbols’ (Robinson 1969a, p. 47).” [Erhardt 2024, p. 10]
@@ 3.5. Platonism and consistency
[T]he mathematician tacitly adopts a realist outlook, imagining herself to investigate an independent realm of mathematical objects … about which she can discern objective truths. This assumption is crucial to the way mathematicians engage with their subject. [Erhardt 2024, pp. 1–2; emphasis added]
Erhardt’s ‘independent realm’ becomes an ‘independent reality’ by page 15. Postulating an ‘independent realm of mathematical objects’ is the gist of a Platonist position. Thus Erhardt postulates that being a Platonist is ‘crucial’ for the working mathematician. Or is it? Sir Michael Atiyah, the recipient of both a Fields Medal and an Abel Prize who was evidently not unsuccessful in “engaging with his subject”, had the following to say concerning independent realms:
@@
[Robinson] does not consider the possibility that even if there are objective answers, they may be beyond our ability to know. [Erhardt 2024, p. 3; emphasis added]
@@
On page 15, Erhardt repeats the unfounded accusation aimed at Robinson, of “deflating mathematics to a game” [Erhardt 2024, p. 15]. There follows a curious paragraph on page 15 in Erhardt concerning Platonism and the problem of consistency:
@@
Robinson does not venture to explain why he takes infinitary mathematics to be consistent, though a compelling answer might have provided a rationale for accepting the finitary results it produces. In the absence of such an explanation, relying upon infinitary mathematics is itself a source of great risk, for there is no assurance that the derived concrete implications are valid. [Erhardt 2024, p. 15; emphasis added]
@@ 3.6. Grand justification wherefore?
[Robinson] neglects to pursue a grand justification for infinitary mathematics by–in lieu of attempting to offer ‘a satisfactory intellectual motivation’–deflating mathematics to a game. [Erhardt 2024, p. 15; emphasis added].
@@
While Robinson repudiates actual infinity, the mathematics of his career revolves around it. This incongruity mirrors the move from (i) to (ii); … Robinson says the same, noting that his position suffers a serious drawback because ‘the gap due to the absence of consistency proofs for the major mathematical theories appears to be inevitable and we have learned to live with it’. [Erhardt 2024, p. 10; emphasis added]
@@
The inevitability of such a gap has been challenged in [Artemov 2024]. Some logicians are skeptical about the philosophical implications of Artemov’s result. Regardless of the outcome of that particular debate, there seems to be little reason to assume that incompleteness is less of a problem for Platonism than for Formalism, and Robinson certainly did not claim such a thing.
@@ 4.1. Meanings of Standardness
“Adding to the irony–and irresolvable tension in Robinson’s work and philosophy–is that, due to its predication on the transfinite, Robinson himself says that ‘the entire notion of standardness must be meaningless […]’ (Robinson 1964/1979b, p. 242).” [Erhardt 2024, note 33; emphasis added]
@@ 4.2. Resolving tensions
(2) On the other hand, to make sense of the question whether there are arbitrarily large numbers $n$ for which $110^{n}+$ 1 is prime does require some theory of arithmetic, and perhaps (if Erhardt’s reading of Quine is right) also an infinite set ${\mathbb{N}}$.
@@ 4.3. Varieties of realism
Consider Cantor’s continuum problem: Is there an intermediate cardinality between the naturals and the reals? If you are a realist about infinite totalities, then the continuum problem is an obvious question to ask. However, ZFC+CH and ZFC+CH [sic] [^17] The second occurrence of ZFC+CH is a typo and should be “ZFC+$\neg$ CH”. are both consistent, which Robinson takes as implying that there must be no universe of sets. A platonist will respond that we have simply not yet found the axioms stating the basic properties of the universe of sets sufficient to decide CH one way or the other. This is a valid objection, but it does not sway Robinson. [Erhardt 2024, note 5; emphasis added]
@@ 4.4. Applicability of mathematics
[Robinson’s] foundational position undermines not only the use of nonstandard analysis, but also Robinson’s considerable corpus of pre-logic contributions [^19] See note 8. to the field in such diverse areas as differential equations and aeronautics [Erhardt 2024, Abstract]
@@ 4.5. Gaifman’s realism
Note that “$\phi$ is provable” and “$T$ is consistent” are respectively $\Sigma^{0}_{1}$ and $\Pi^{0}_{1}$ sentences (when encoded in PA). Gaifman apparently holds that one commits oneself to realism in truth value (in relation to the natural numbers) when one uses the terms provable and consistent in metamathematics. He seems to find it philosophically unacceptable to assume that metamathematical statements about the provability of a sentence or the consistency of a theory could be undecidable. Meanwhile, for a formalist practicing finitistic metamathematics in Primitive Recursive Arithmetic (PRA) such a posture toward undecidability is a natural consequence. The “finitistic semantics” of PRA is naturally incomplete; see [Tait 1981]. A $\Pi^{0}_{1}$ sentence is ‘true’ if it can be proved by induction; it is ‘false’ if one can provide a counterexample. A $\Sigma^{0}_{1}$ sentence is ‘true’ if one can provide an example; it is ‘false’ if one can provide an inductive proof of the corresponding $\Pi^{0}_{1}$ sentence expressing its negation; see further in [Simpson 2009].
@@ 4.6. Knowledge-transcendent truth?
[…]
@@
Gödel’s second incompleteness theorem as applied to PA is a classical arithmetic independence result. Gaifman appears to claim that such a result would “reinforce realism.” Accordingly, the independence of CON(PA) of PA would reinforce realism. One wonders what Gaifman would make of Artemov’s recent proof of CONS(PA) within PA.222222[Artemov 2024] proposed an alternative formalisation ConS(PA) of consistency following Hilbert. Giaquinto quotes Robinson in [Robinson 1965, p. 239] to the effect that “I cannot see at this time how a form of reasoning which attempts to escape the consequences of Gödel’s second theorem (such as Gentzen’s consistency proof or any other consistency proof for Arithmetic) can remain strictly finitistic and hence interpreted” and claims that “the formalist foundational programme is irredeemable” [Giaquinto 1983, p. 129]. Artemov’s result provides a possible way of escaping the consequences of Gödel’s second theorem, and therefore possibly undermines Giaquinto’s rash conclusion. Similar remarks apply to Giaquinto’s surprising claim that “the resuscitation of formalism by Robinson and Cohen (following Cohen’s proof of the independence of the Continuum Hypothesis) is untenable and barren” [Giaquinto 1983, p. 120]. If even independence reinforces realism, lack of independence (in the sense of Artemov’s result) would seem to also reinforce realism. This would suggest that Gaifman’s variety of realism is unfalsifiable.

### v3

@@ 1. Formalism and finitism
Abraham Robinson’s intellectual profile continues to inspire passions fifty years after his passing. A number of publications over the past few decades have analyzed his philosophical position, including [Benis-Sinaceur 1988, p. 603], [Gaifman 2012], [Sherry and Katz 2012], [Sanders 2020], [Weir 2024], and most recently [Erhardt 2025]. A majority of commentators take it for granted that Robinson’s mature position is best described as mathematical Formalism. However, Erhardt reads Robinson as a finitist (see Section 1.2 for a discussion of distinct meanings of the term) who rejects infinite totalities.
@@ 1.1. Meaning and reference
In Formalism 64, Robinson’s most fully developed philosophical paper, he writes (Robinson 1964/1979b, pp. 230–231, italics in the original), “[…] the notion of a particular class of five elements, e.g. of five particular chairs, presents itself to my mind as clearly as the notion of a single individual […]. By contrast, I feel quite unable to grasp the idea of an actual infinite totality. To me there appears to exist an unbridgeable gulf between sets or structures of one, or two, or five elements, on one hand, and infinite structures on the other hand […]” [^2] Robinson as quoted in [Erhardt 2025, p. 431]; emphasis on “unable to grasp” added.
@@ 1.2. Gaifman’s reading of Robinson
Erhardt thanks Gaifman for encouraging him to read Robinson’s text Formalism 64 [Erhardt 2025, p. 446]. Gaifman may have been the original source of a characterisation of Robinson as a finitist. Indeed, Gaifman claims in 2012:
@@ 1.3. Standard model
There is, however, another alternative: more strongly committing to finitism. This is the path taken by various constructivists, who rather than play the uninterpretable game of symbols choose, in a philosophically principled manner, to adopt a weaker form of mathematics. [Erhardt 2025, p. 444; emphasis added]
@@ 1.4. Robinson’s passage in context
As noted in Section 1.1, Robinson clarified his adjective ‘meaningless’ in his 1975 text where he stated that what he has in mind is an absence of a reference; namely, the term infinite totality does not refer to anything in either the real or a Platonic realm. Already in Formalism 64 he emphasized that he rejected the idea that infinite totalities exist “either really or ideally” [Robinson 1965, p. 230]. Here ‘really’ refers to the physical realm, and ‘ideally’ to a Platonic realm, as he states explicitly in the passage quoted above. It emerges that Erhardt’s presentation of Robinson’s position, as when Erhardt claims that it “places Robinson at odds with typical mathematical practice” [Erhardt 2025, p. 432], conflates the generic meaning of terms such as to grasp and meaningless, with the precise technical meaning attributed to such terms by Robinson.
@@ 1.5. Fermat’s Last Theorem: is the proof legitimate?
Though Fermat’s Last Theorem is a general result proven by illegitimate methods and that purports to say something about all natural numbers—constituting an illicit reference to actual infinity—one can derive from it a potentially infinite number of legitimate, material claims. Yet the fact that the proof is illegitimate on Robinson’s view is crucial to our appraisal of his view. [Erhardt 2025, p. 441; emphasis added]
@@
In opposition to nominalists (and a fortiori, fictionalists), Robinson is a devout realist with regard to each individual natural number. This is evidenced by his commitment to potential infinity, …777[Erhardt 2025, p. 438, note 22]; emphasis on “devout realist” added.
@@
With regard to the issue of nominalism and fictionalism mentioned by Erhardt above, we note a pecularity of Erhardt’s approach. Erhardt is keen to cast Robinson as an instrumentalist, rather than one who regards infinite classes as useful fictions. Erhardt dismisses the suggestion that Robinson was a fictionalist, in spite of acknowledging that Leibniz, an avowed fictionalist, “had a profound effect” on Robinson [Erhardt 2025, p. 437 and note 22]. The rationale for such a dismissal is most peculiar: Erhardt considers that it would be anachronistic for Robinson, in 1964, to see himself as a fictionalist since Field’s ‘introduction of fictionalism’ occurred only in 1980 (see [Field 1980]). Erhardt is right to note that Field’s fictionalism, a form of nominalism, conflicts with Robinson’s repeated rejection of nominalism. But such a conflict arises only if fictionalism is assumed to be necessarily a nominalist proposal. Leibniz’s fictionalism offers an alternative to nominalism. Leibniz treated infinitesimals, along with negative and imaginary numbers, as well-founded fictions; see [Sherry and Katz 2012]. Although he sometimes suggested that infinitesimals are eliminable in the manner of Archimedean exhaustion arguments, he made no such suggestion for negatives and imaginaries. In all three cases, it is the contribution to systematicity that establishes the mantle ‘well-founded fiction’, rather than any sort of nominalist reduction.
@@ 1.6. Symptomatic keyword list and imaginary tensions
The foundational position he inherited from David Hilbert undermines not only the use of nonstandard analysis, but also Robinson’s considerable corpus of pre-logic contributions [^8] Erhardt’s wording here constitutes a historical error. He makes it appear as though Robinson’s work in aeronautics, such as his 1956 book Wing Theory [Robinson and Laurmann 1956], preceded his work in logic. This is incorrect. For example, Robinson’s dissertation The metamathematics of algebraic systems [Robinson 1949], advised by Paul Dienes, dates from 1949; see https://www.genealogy.math.ndsu.nodak.edu/id.php?id=15886. See also Dauben [Dauben 1995, p. 157]. to the field in such diverse areas as differential equations and aeronautics. This tension emerges from Robinson’s disbelief in the existence of infinite totalities … [Erhardt 2025, Astract; emphasis added]
@@ 2. Twin prime conjecture and types of infinite totalities
[E]ven basic conjectures of number theory, such as the twin prime conjecture, presuppose the existence of infinite totalities. [Erhardt 2025, p. 431]
However, such a claim involves a conflation of distinct meanings associated with the term infinite totality; namely, of different levels of language or theory. The twin prime conjecture (TPC) can be formulated in Peano Arithmetic (PA). This can be done, for example, by the formula on page 442, line 4 in [Erhardt 2025], which we will express as follows:
@@ 3.1. A logical fallacy
While Robinson has no difficulty grasping the possibility of arbitrarily large finite sets, which motivates his acceptance of the natural numbers taken individually, he reports that it is simply a psychological fact that he cannot grasp infinitary objects, obviating the possibility of finding common ground with disputants.101010[Erhardt 2025, p. 431]; emphasis on ‘grasping’ in the original; emphasis on ‘reports’ and ‘psychological’ added.
@@ 3.2. Robinson’s letter to Gödel
Erhardt’s attribution of Alice’s position to Robinson constitutes a logical fallacy [^13] Robinson is not the only mathematician whose position is misrepresented by Erhardt. Erhardt claims that Edward Nelson “believe[d] there is a largest number” [Erhardt 2025, p. 433]. To support his claim, Erhardt cites four texts by Nelson. We examined all four, including Nelson’s publication [Nelson 2011], but found no evidence of such a belief as claimed by Erhardt. Erhardt’s claim in [Erhardt 2025, p. 433 note 7] that Nelson “believed Peano Arithmetic was inconsistent” is similarly inaccurate. Rather, Nelson suspected that PA was inconsistent. When he thought he had a proof he obviously thought it was inconsistent, but when Tao found an error in the proof, Nelson immediately accepted it. and does not inspire confidence in his analysis.
@@ 3.3. Verbal excesses from Kreisel to Erhardt
Robinson has already admitted that such totalities are meaningless, and without an argument defending the use of these languages, formalism becomes an inconsistent project [Erhardt 2025, p. 435; emphasis added].
@@ 3.4. Whose game was it anyway?
“He explains non-finitary mathematics as a collection of ‘uninterpretable games with symbols’ (Robinson 1969a, p. 47).” [Erhardt 2025, p. 439]
@@ 3.5. Platonism and consistency
[T]he mathematician tacitly adopts a realist outlook, imagining herself to investigate an independent realm of mathematical objects … about which she can discern objective truths. This assumption is crucial to the way mathematicians engage with their subject. [Erhardt 2025, pp. 429–430; emphasis added]
Erhardt’s ‘independent realm’ becomes an ‘independent reality’ by page 444 Postulating an ‘independent realm of mathematical objects’ is the gist of a Platonist position. Thus Erhardt postulates that being a Platonist is ‘crucial’ for the working mathematician. Or is it? Sir Michael Atiyah, the recipient of both a Fields Medal and an Abel Prize who was evidently not unsuccessful in “engaging with his subject”, had the following to say concerning independent realms:
@@
[Robinson] does not consider the possibility that even if there are objective answers, they may be beyond our ability to know. [Erhardt 2025, p. 432; emphasis added]
@@
On page 444, Erhardt repeats the unfounded accusation aimed at Robinson, of “deflating mathematics to a game” [Erhardt 2025, p. 444]. There follows a curious paragraph on page 15 in Erhardt concerning Platonism and the problem of consistency:
@@
Robinson does not venture to explain why he takes infinitary mathematics to be consistent, though a compelling answer might have provided a rationale for accepting the finitary results it produces. In the absence of such an explanation, relying upon infinitary mathematics is itself a source of great risk, for there is no assurance that the derived concrete implications are valid. [Erhardt 2025, p. 444; emphasis added]
@@ 3.6. Grand justification wherefore?
[Robinson] neglects to pursue a grand justification for infinitary mathematics by–in lieu of attempting to offer ‘a satisfactory intellectual motivation’–deflating mathematics to a game. [Erhardt 2025, p. 444; emphasis added].
@@
While Robinson repudiates actual infinity, the mathematics of his career revolves around it. This incongruity mirrors the move from (i) to (ii); … Robinson says the same, noting that his position suffers a serious drawback because ‘the gap due to the absence of consistency proofs for the major mathematical theories appears to be inevitable and we have learned to live with it’. [Erhardt 2025, p. 439; emphasis added]
@@
The inevitability of such a gap has been challenged in [Artemov 2025]. Some logicians are skeptical about the philosophical implications of Artemov’s result. Regardless of the outcome of that particular debate, there seems to be little reason to assume that incompleteness is less of a problem for Platonism than for Formalism, and Robinson certainly did not claim such a thing.
@@ 4.1. Meanings of Standardness
“Adding to the irony–and irresolvable tension in Robinson’s work and philosophy–is that, due to its predication on the transfinite, Robinson himself says that ‘the entire notion of standardness must be meaningless […]’ (Robinson 1964/1979b, p. 242).” [Erhardt 2025, p. 442 note 33; emphasis added]
@@ 4.2. Resolving tensions
(2) On the other hand, to make sense of the question whether there are arbitrarily large numbers $n$ for which $10^{n}+1$ is prime does require some theory of arithmetic, and perhaps (if Erhardt’s reading of Quine is right) also an infinite set $\mathbb{N}$}.
@@ 4.3. Varieties of realism
[…]
@@ 4.4. Applicability of mathematics
[Robinson’s] foundational position undermines not only the use of nonstandard analysis, but also Robinson’s considerable corpus of pre-logic contributions [^19] See note 8. to the field in such diverse areas as differential equations and aeronautics [Erhardt 2025, Abstract]
@@ 4.5. Gaifman’s realism
Note that $\phi$ is provable” and “$T$ is consistent” are respectively $\Sigma^{0}_{1}$ and $\Pi^{0}_{1}$ sentences (when encoded in PA). Gaifman apparently holds that one commits oneself to realism in truth value (in relation to the natural numbers) when one uses the terms provable and consistent in metamathematics. He seems to find it philosophically unacceptable to assume that metamathematical statements about the provability of a sentence or the consistency of a theory could be undecidable. Meanwhile, for a formalist practicing finitistic metamathematics in Primitive Recursive Arithmetic (PRA) such a posture toward undecidability is a natural consequence. The “finitistic semantics” of PRA is naturally incomplete; see [Tait 1981]. A $\Pi^{0}_{1}$ sentence is ‘true’ if it can be proved by induction; it is ‘false’ if one can provide a counterexample. A $\Sigma^{0}_{1}$ sentence is ‘true’ if one can provide an example; it is ‘false’ if one can provide an inductive proof of the corresponding $\Pi^{0}_{1}$ sentence expressing its negation; see further in [Simpson 2009].
@@ 4.6. Knowledge-transcendent truth?
realism $\Rightarrow\text{\small independence results indicate that truths outpace our capacities}$
$\Rightarrow\text{\small there exist transcendent truths}$
$\Rightarrow\text{\small realism}.$
@@
Gödel’s second incompleteness theorem as applied to PA is a classical arithmetic independence result. Gaifman appears to claim that such a result would “reinforce realism.” Accordingly, the independence of CON(PA) of PA would reinforce realism. One wonders what Gaifman would make of Artemov’s recent proof of CONS(PA) within PA.222222[Artemov 2025] proposed an alternative formalisation ConS(PA) of consistency following Hilbert. Giaquinto quotes Robinson in [Robinson 1965, p. 239] to the effect that “I cannot see at this time how a form of reasoning which attempts to escape the consequences of Gödel’s second theorem (such as Gentzen’s consistency proof or any other consistency proof for Arithmetic) can remain strictly finitistic and hence interpreted” and claims that “the formalist foundational programme is irredeemable” [Giaquinto 1983, p. 129]. Artemov’s result provides a possible way of escaping the consequences of Gödel’s second theorem, and therefore possibly undermines Giaquinto’s rash conclusion. Similar remarks apply to Giaquinto’s surprising claim that “the resuscitation of formalism by Robinson and Cohen (following Cohen’s proof of the independence of the Continuum Hypothesis) is untenable and barren” [Giaquinto 1983, p. 120]. If even independence reinforces realism, lack of independence (in the sense of Artemov’s result) would seem to also reinforce realism. This would suggest that Gaifman’s variety of realism is unfalsifiable.